Find the value or values of y in the quadratic equation y^2+4y+4=7
Answer by checkley71(8403)

Y~2=4Y=4=7 OR (Y+2)(Y+2)=7 OR (Y+2)~2=7 OR TAKING THE SQUARE ROOT OF BOTH SIDES
Y+2=SQRT7 OR Y=-2+SQRT7.